OpenStudy (anonymous):

Simplify.\[\Large\tt\frac{ n }{ 2-n }-\frac{ 2 }{ n-2 }\]

OpenStudy (anonymous):

I need to get them to have the same denominator. Can you help?

OpenStudy (anonymous):

You can rewrite the second as\[\frac{ -2 }{ 2-n }\]

OpenStudy (anonymous):

How do you get to that? Do you multiple by a negative?

OpenStudy (anonymous):

You can multiply any number by 1 and still get the same number (multiplicative identity) So using that you can multiply a fraction by \[\frac{ -1 }{ -1 }\]. the rest is reordering and combining

OpenStudy (anonymous):

Oh, okay! I understand now. Thanks...

OpenStudy (anonymous):

So the final answer here after combining would be \[\frac{ n+2 }{ 2-n }\]

OpenStudy (anonymous):

Got it! Thank you very much! :D